# Past, Present, and Future of Bug Tracking in the Generative AI Era

### v1

@@ Abstract.
Traditional bug tracking systems rely heavily on manual reporting, reproduction, triaging, and resolution. These tasks are performed by different stakeholders such as end users, customer support, developers, and testers. This division of responsibilities requires significant coordination in addition to the considerable human time and effort. This reliance widens the communication gap between non-technical end users and technical developers, slowing the process from bug discovery to resolution and deployment. Moreover, current solutions tend to be highly asynchronous; end users often wait hours, days, or even weeks before receiving an initial response, further delaying fixes and contributing to frustration.
In this paper, we examine the evolution of bug tracking practices, moving from early paper-based and manual reporting methods to today’s web-based and software as a service (SaaS) platforms that dominate modern software development. Building on this trajectory, we outline our vision for the future: an AI-powered bug tracking framework that augments existing systems with intelligent, large language model (LLM) driven automation.
Our approach tackles two main challenges: reducing the time to resolution (TTR) and minimizing coordination overhead by bridging the gap between end-users and developers. In the proposed framework, end users report bugs in natural language. AI-driven agents refine these reports, attempt to reproduce them, and request missing details, if any. Bug reports are classified according to their type. Invalid bugs are automatically classified and either dismissed or resolved via no-code fixes suggested by agents. Valid bugs, once confirmed, are localized and assigned to developers, who review LLM-generated patches and verify their correctness. Verified patches are deployed using continuous integration and continuous development (CI/CD) pipelines, which are already generated and maintained by the LLM agents.
This paper presents our vision for an AI-powered bug tracking framework, highlighting the challenges and opportunities of integrating LLMs into bug tracking. By positioning our framework as both a forward-looking design and a practical augmentation for existing bug tracking tools, we try to illustrate how automation not only decreases response times and TTR but also transforms the overall software maintenance process, paving the way for a more efficient, collaborative, and user-centric future.
@@
journal: TOSEMjournalvolume: 1journalnumber: 1article: 1publicationmonth: 10copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXccs: Software and its engineering Software maintenance toolsccs: Artificial intelligence Intelligent agents
@@ 1. Introduction
Over time, bug tracking has evolved into an essential backbone of collaborative software development, becoming integrated into broader software engineering workflows. Modern bug tracking practices emphasize traceability and coordination among diverse stakeholders (Anvik et al., 2006), enabling teams to track the entire lifecycle of a bug, from discovery and reporting to resolution and verification. Bug tracking repositories now serve not only as records of defects but also as valuable sources of knowledge, informing project management decisions, supporting quality assurance, and guiding future development (Zimmermann et al., 2010). Despite these advancements, the fundamental manual nature of bug tracking has remained largely unchanged. Thus, problems might arise in the names of communication problems, ambiguous bug reports (Zimmermann et al., 2010), and the need for extensive human intervention (Anbalagan and Vouk, 2009).
@@
As the volume of reported bugs grows with the scale of software projects, the importance of efficient and reliable bug tracking systems has become more pronounced, solidifying their role as a critical enabler of software quality assurance (Hassan, 2009; Guo et al., 2010). Widely used bug tracking tools such as Jira [^1] https://www.atlassian.com/software/jira, Bugzilla [^2] https://www.bugzilla.org/, and GitHub Issues [^3] https://github.com/features/issues have played central roles in software engineering. Jira, released in 2002, has become a dominant system in enterprise settings for handling bugs, tasks, and project management workflows. Bugzilla, publicly available since 1998, has been used in many open source projects as one of the canonical issue tracking platforms, particularly for its flexibility and openness. GitHub Issues has also emerged as one of the most widely used bug tracking systems in open source communities. These tools are used widely across open source communities and proprietary organizations, often acting as the canonical system for bug reporting, triage, assignment, status tracking, historical analysis, and the communication channel for the stakeholders.
@@
However, despite their maturity and widespread adoption, these systems have notable deficiencies. Jira and Bugzilla often impose overhead by requiring detailed workflow configuration and many mandatory fields, which can slow down bug reporting and triage. For example, Jira users have complained about performance and usability issues, like the system slowing down when workflows get too complex [^4] https://jira.atlassian.com/browse/JRASERVER-61704. In contrast, GitHub Issues has been criticized for its unstructured design. Empirical studies show that many issues are created without sufficient context, lacking reproducible steps or environment details (Shah et al., 2025). Software maintainers often cannot, or choose not to, enforce strict templates or workflows in many repositories. For example, a study on large GitHub projects (Sülün et al., 2024) found that the ontology of issue templates and required metadata is either minimal or inconsistently used, leading to ambiguity in many reports. These deficiencies result in several inefficiencies, including duplicate reports, slower triage and resolution, inconsistent issue quality, and higher manual efforts.
@@
The recent rise of large language models (LLMs) such as Gemini (Comanici et al., 2025), GPT (Achiam et al., 2023), and Llama (Touvron et al., 2023) has transformed natural language processing by enabling machines to generate, understand, and reason over text at unprecedented levels of fluency and accuracy. Leveraging massive pre-training on diverse text corpora, these models demonstrate remarkable zero-shot and few-shot capabilities, generalizing across a wide variety of tasks without requiring task-specific training. Beyond research benchmarks, LLMs are increasingly deployed in real-world applications such as conversational assistants, code generation, and knowledge management, fundamentally changing the way humans interact with software systems. Their ability to interpret unstructured inputs, extract structured information, and produce contextually relevant outputs suggests strong potential for integration into complex workflows that usually exist in software engineering. GitHub Copilot [^5] https://github.com/features/copilot, for instance, has already shown how LLMs can be used in programming tasks by providing real-time code suggestions, signaling a shift toward AI-augmented software development (Vaithilingam et al., 2022).
@@
Beyond standalone LLMs, researchers are increasingly focusing on LLM agents, autonomous or semi-autonomous systems that can understand their environment, plan tasks, and carry out actions to achieve long-term goals (Wang et al., 2024c). These agents can be combined into multi-agent systems, where multiple LLMs collaborate through structured communication, task specialization, and coordination protocols (Hong et al., 2024; Park et al., 2023). Such orchestration allows agents to complement each other, mitigate individual weaknesses, and scale to solve more sophisticated problems than a single model could handle in isolation. Frameworks such as AutoGPT [^6] https://github.com/Significant-Gravitas/AutoGPT and MetaGPT [^7] https://github.com/FoundationAgents/MetaGPT exemplify this direction, where agents assume distinct roles (e.g., planner, coder, tester) and cooperate to achieve complex objectives. The emergence of these paradigms highlights the feasibility of constructing collaborative ecosystems of LLMs that can continuously refine, validate, and operationalize knowledge in dynamic domains.
@@
Although there are some studies on this topic, the integration of LLMs into bug tracking is still relatively underexplored. Nonetheless, it remains an important direction because traditional methods often face communication problems [^8] https://github.com/huggingface/transformers/issues/39803 , slow issue resolution [^9] https://github.com/huggingface/transformers/issues/17283 , and high manual effort [^10] https://github.com/huggingface/transformers/issues/40136 , as shown in the footnoted examples. While recent works have applied LLMs to some tasks in bug tracking, such as bug localization (Xia et al., 2023; Lyu et al., 2024) and program repair (Zhou et al., 2024; Nong et al., 2024), others have explored their use in broader software engineering activities (Hou et al., 2024; Hassan et al., 2024; Patil and Jadon, 2024). Yet, the integration of LLM agents into the end-to-end bug tracking lifecycle, from bug report generation to bug verification and patch deployment, has not been systematically studied. Moreover, the potential of collaborative LLM agents to act as automated triagers, verifiers, or explainers of bug reports has yet to be fully realized. This gap presents a promising research opportunity to reimagine bug tracking systems by embedding LLMs and multi-agent collaboration into their core, thereby reducing manual effort, improving report quality, and accelerating defect resolution.
@@
By automating repetitive and time-consuming tasks and bridging the end-users, who are usually non-technical, and the technical developers, our vision significantly reduces the time and human effort required for bug triaging and resolution. It also addresses communication issues [^11] https://github.com/huggingface/transformers/issues/39803 that often arise during bug tracking. Through automation, we aim to minimize human involvement in repetitive tasks [^12] https://github.com/huggingface/transformers/issues/40136 that typically happen during bug tracking while maintaining high accuracy and efficiency in bug resolution. We discuss the potential challenges and long-term research opportunities in integrating AI-driven automation into software maintenance workflows, specifically in bug tracking. We envision a future where AI and human collaboration, within a human-in-the-loop (HIL) system, redefines software maintenance.
@@
2. Historical Evolution and Current System of Bug Tracking
This section provides an analysis of the bug-tracking process, from its foundational principles to its modern implementations. Section 2.1 reviews the evolution of bug tracking systems, showing how early manual approaches provided the foundation for today’s advanced platforms. Section 2.2 describes the architecture and workflow of the current bug tracking systems, detailing their key components and how they function within the software development lifecycle. Finally, Section 2.3 critically evaluates the challenges of these current systems, identifying the challenges and inefficiencies.
@@ 2.1.2. Pre-Internet Era (1970s-1980s)
During this period, communication began to evolve beyond face-to-face meetings. Email systems [^13] https://en.wikipedia.org/wiki/ALL-IN-1141414https://en.wikipedia.org/wiki/IBM_OfficeVision, simple databases like dBase [^15] https://www.dbase.com, and early file-sharing networks emerged. Bugs were often reported by users via email or phone calls to customer support. The software itself was shipped on floppy disks, and users would sometimes mail back faulty disks along with registration forms. When a customer support representative received a bug report, they would write the reproduction steps into a text file. This file would then be passed to a developer who would manually try to reproduce the bug. Fixing a bug involved manual code changes, which were then shared with the customer via a new floppy disk or, if available, a local network. Verification was done through manual testing with a documented checklist. The process was very slow and inefficient. While remote communication was a new development, collaboration was still low, often leading to a high TTR.
@@ 2.1.3. Internet Era (1980s-1990s)
This era saw the first dedicated bug-tracking systems. Tools like GNATS (GNU Bug Tracking System) [^16] https://www.gnu.org/software/gnats/, Debuggs [^17] https://en.wikipedia.org/wiki/Debbugs, and CMVC (Yu, 1994) were introduced. GNATS, for instance, was an early, open source bug tracking system that used text-based files and email for communication, providing a structured interface for logging and tracking bugs. Debbugs was the software powering the Debian project’s issue tracking system. It did not have any form of web interface to edit bug reports, all modification was done through email. These systems allowed for detailed reproduction steps to be logged and were accessible to the whole team. When a bug was fixed, the code was committed to a version control system (VCS), and the bug report was updated with details about the fix.
@@
To manage the bug backlog, Netscape [^18] https://isp.netscape.com/ team leaders held weekly bug meetings, or ”bug councels”. As the project neared its release, these meetings increased in frequency to once or twice a day to prioritize and make decisions on fixes. These meetings were typically led by a project or release manager and included representatives from development, QA, and marketing teams. Netscape used a five-level classification system: critical, major, normal, minor, and trivial. Bugs were classified and entered into the database by the engineers who found them, typically from the QA team. The marketing team also maintained a ”top 10 bugs” list based on customer feedback, which was passed to QA (Cusumano and Yoffie, 1999). By the mid-to-late 1990s, centralized client-server tools began to surface, offering more structured workflows.
@@ 2.1.4. Web-Based Era (2000s)
This decade saw a major shift to web-based platforms. Tools like Bugzilla, MantisBT [^19] https://mantisbt.org/, Trac [^20] https://trac.edgewall.org/, and early versions of Jira emerged. Bugzilla became a landmark open source system. Bugzilla introduced key abstractions still in use today, including structured fields (e.g., severity, priority, component), status transitions, and user roles. It provided a web-based interface and notification mechanisms, enabling better coordination among developers and testers (Mockus et al., 2002). These systems treated bug reports as work items in a sprint, integrating with the emerging agile methodologies. The web UI made it much easier for everyone to access and track bugs. The expansion of community-contributed plug-ins and themes also reflected the diverse needs of development teams, from small startups to large open source foundations (Bertram et al., 2010). Collaboration improved with shared dashboards and email notifications, though bottlenecks were still common due to a lack of deep integration with development tools (Eren et al., 2023).
2.1.5. Software as a service (SaaS), DevOps and Automation Era (2010s-2022)
In this era, platforms like GitHub Issues, GitLab Issues [^21] https://docs.gitlab.com/user/project/issues/, YouTrack [^22] https://www.jetbrains.com/youtrack/, and Azure DevOps [^23] https://azure.microsoft.com/en-us/products/devops became the standard. Bug tracking became fully integrated into the development lifecycle. GitHub Issues, for example, is part of a larger ecosystem that includes version control and continuous integration. Fixes were no longer just committed; they became a part of a Continuous Integration/Continuous Deployment (CI/CD) pipeline, where code changes automatically trigger tests and deployments.
The growing popularity of issue-driven development, where code commits are directly tied to issue identifiers, further solidified the role of bug tracking systems as central elements of the software delivery process. Developers began to view bug tracking systems not merely as databases of bugs, but as vital components of planning, communication, and accountability (Gousios et al., 2014; Bird et al., 2009).

### v2

@@ Abstract.
Traditional bug tracking systems rely heavily on manual reporting, reproduction, classification, and resolution. These tasks are performed by various stakeholders, including end users, customer support, developers, and testers. This division of responsibilities requires significant coordination in addition to the considerable human time and effort. This reliance widens the communication gap between non-technical end users and developers, who are more technically inclined, slowing the process from bug discovery to resolution and deployment. Moreover, current solutions tend to be highly asynchronous; end users often wait hours, days, or even weeks before receiving an initial response, further delaying fixes and contributing to frustration.
In this paper, we examine the evolution of bug tracking practices, moving from early paper-based and manual reporting methods to today’s web-based platforms that dominate modern software development. Building on this trajectory, we outline our vision for the future: an AI-powered bug tracking framework that augments existing systems with intelligent, large language model (LLM) and agent-driven automation, and we report on early adaptations of key components of this framework that provide initial empirical grounding for its feasibility.
Our approach tackles two main challenges: reducing the time to resolution and minimizing coordination overhead by bridging the gap between end-users and developers. In the proposed framework, end users report bugs in natural language. AI-driven agents refine these reports, attempt to reproduce them, and request any missing details. Bug reports are classified according to their severity, priority, and type. Valid and invalid bugs are automatically determined, and the invalid bug reports are either dismissed or resolved via no-code fixes suggested by agents. Valid bugs, once confirmed, are localized and assigned to developers, who review LLM-generated patches and verify their correctness. Verified patches are deployed using continuous integration and continuous development (CI/CD) pipelines.
This paper articulates our vision for an AI-powered bug tracking framework and examines the challenges and opportunities of integrating LLMs into bug tracking workflows. We position the framework as a forward-looking design that can also augment today’s issue trackers, illustrating how automation can reduce turnaround time while reshaping software maintenance into a more efficient, collaborative, and user-centric process.
@@
journal: TOSEMjournalvolume: 1journalnumber: 1article: 1publicationmonth: 1copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXccs: Software and its engineering Software maintenance toolsccs: Artificial intelligence Intelligent agents
@@ 1. Introduction
Over time, bug tracking has evolved into an essential backbone of collaborative software development, becoming integrated into broader software engineering workflows. Modern bug tracking practices emphasize traceability and coordination among diverse stakeholders (Anvik et al., 2006), enabling teams to track the entire lifecycle of a bug, from discovery and reporting to resolution and verification. Bug tracking repositories now serve not only as records of defects but also as valuable sources of knowledge, informing project management decisions, supporting quality assurance, and guiding future development (Zimmermann et al., 2010). Despite these advancements, the fundamental manual nature of bug tracking has remained largely unchanged. Consequently, the process is hindered by communication gaps, ambiguous reports (Zimmermann et al., 2010), and high reliance on human intervention (Anbalagan and Vouk, 2009).
@@
The correlation between increasing project scale and bug frequency underscores the necessity of robust bug tracking systems as a cornerstone of software quality assurance (Hassan, 2009; Guo et al., 2010). Widely used bug tracking tools such as Jira [^1] https://www.atlassian.com/software/jira, Bugzilla [^2] https://www.bugzilla.org/, and GitHub Issues [^3] https://github.com/features/issues have played central roles in software engineering. Jira, released in 2002, has become a dominant system in enterprise settings for handling bugs, tasks, and project management workflows. Bugzilla, publicly available since 1998, has been used in many open source projects as one of the canonical issue tracking platforms, particularly for its flexibility and openness. GitHub Issues has also emerged as one of the most widely used bug tracking systems in open source communities. These tools are used widely across open source communities and proprietary organizations, often acting as the canonical system for bug reporting, triage, assignment, status tracking, historical analysis, and the communication channel for the stakeholders.
@@
However, despite their maturity and widespread adoption, these systems have notable deficiencies. These systems might be thought of as serving as an interface for database operations. Moreover, Jira and Bugzilla often impose overhead by requiring detailed workflow configuration and many mandatory fields, which can slow down bug reporting and triage. For example, Jira users have complained about performance and usability issues, like the system slowing down when workflows get too complex [^4] https://jira.atlassian.com/browse/JRASERVER-61704. In contrast, GitHub Issues has been criticized for its unstructured design. Empirical studies show that many issues are created without sufficient context, lacking reproducible steps or environment details (Shah et al., 2025). Software maintainers often cannot, or choose not to, enforce strict templates or workflows in many repositories. For example, a study on large GitHub projects (Sülün et al., 2024) found that the schema of issue templates and required metadata is either minimal or inconsistently used, leading to ambiguity in many reports. These deficiencies result in several inefficiencies, including duplicate reports, slower triage and resolution, inconsistent issue quality, and higher manual efforts.
@@
The recent rise of large language models (LLMs) such as Gemini (Comanici et al., 2025), GPT (Achiam et al., 2023), and Llama (Touvron et al., 2023) has transformed natural language processing (NLP) by enabling machines to generate, understand, and reason over text at unprecedented levels of fluency and accuracy. Leveraging massive pre-training on diverse text corpora, these models demonstrate remarkable zero-shot and few-shot capabilities, generalizing across a wide variety of tasks without requiring task-specific training. Beyond research benchmarks, LLMs are increasingly deployed in real-world applications such as conversational assistants, code generation, and knowledge management, fundamentally changing the way humans interact with software systems. Their ability to interpret unstructured inputs, extract structured information, and produce contextually relevant outputs suggests strong potential for integration into complex workflows that usually exist in software engineering. GitHub Copilot [^5] https://github.com/features/copilot, for instance, has already shown how LLMs can be used in programming tasks by providing real-time code suggestions, signaling a shift toward AI-augmented software development (Vaithilingam et al., 2022).
@@
Beyond standalone LLMs, researchers are increasingly focusing on LLM agents, autonomous or semi-autonomous systems that can understand their environment, plan tasks, and carry out actions to achieve long-term goals (Wang et al., 2024c). These agents can be combined into multi-agent systems, where multiple LLM agents collaborate through structured communication, task specialization, and coordination protocols (Hong et al., 2024; Park et al., 2023). Such orchestration allows agents to complement each other, mitigate individual weaknesses, and scale to solve more complex problems than a single model could handle in isolation. Frameworks such as AutoGPT [^6] https://github.com/Significant-Gravitas/AutoGPT and MetaGPT [^7] https://github.com/FoundationAgents/MetaGPT exemplify this direction, where agents assume distinct roles (e.g., planner, coder, tester) and cooperate to achieve complex objectives. The emergence of these paradigms highlights the feasibility of constructing collaborative ecosystems of LLMs that can continuously refine, validate, and operationalize knowledge in dynamic domains.
@@
Integrating LLMs into bug tracking is underexplored but vital, addressing the communication issues [^8] https://github.com/microsoft/vscode/issues/519, resolution delays [^9] https://github.com/kubernetes/kubernetes/issues/117706, and manual overhead [^10] https://github.com/pytorch/pytorch/issues/77764 typical of traditional methods. While recent works have applied LLMs to some tasks in bug tracking, such as bug localization (Xia et al., 2023; Lyu et al., 2024) and automated program repair (Zhou et al., 2024; Nong et al., 2024), others have explored their use in broader software engineering activities (Hou et al., 2024; Hassan et al., 2024; Patil and Jadon, 2024). However, the integration of LLM agents into the whole end-to-end bug tracking lifecycle has not been systematically studied. Moreover, the potential of collaborative LLM agents to act as automated triagers, verifiers, or explainers of bug reports has yet to be fully realized. This gap presents a promising research opportunity to reimagine bug tracking systems by embedding LLMs and multi-agent collaboration into their core, thereby reducing manual effort, improving report quality, and accelerating defect resolution.
@@
By automating repetitive and time-consuming tasks and bridging the gap between end-users, who are usually non-technical, and developers, who form the technical side, our vision aims to reduce the time and human effort required for bug reporting and resolution. It also addresses communication issues [^11] https://github.com/microsoft/vscode/issues/519 that often arise during bug tracking. Through automation, we aim to minimize human involvement in repetitive tasks [^12] https://github.com/pytorch/pytorch/issues/77764 that typically happen during bug tracking while maintaining high accuracy and efficiency in bug resolution. We discuss the potential challenges and long-term research opportunities in integrating AI-driven automation into software maintenance workflows, specifically in bug tracking. We envision a future where AI and human collaboration, within a human-in-the-loop (HIL) system, redefines software maintenance.
@@
2. Historical Evolution and Traditional System of Bug Tracking
This section provides an analysis of the bug tracking process, from its foundational principles to its modern implementations. Section 2.1 reviews the evolution of bug tracking systems, showing how early manual approaches provided the foundation for today’s advanced platforms. Section 2.2 describes the workflow of the traditional bug tracking systems, detailing their key components and how they function within the software development lifecycle. Finally, Section 2.3 critically evaluates the challenges of these traditional systems, identifying the challenges and inefficiencies.
@@ 2.1.2. Pre-Internet Era (1970s-1980s)
During this period, communication began to evolve beyond face-to-face meetings. Email systems [^13] https://en.wikipedia.org/wiki/ALL-IN-1141414https://en.wikipedia.org/wiki/IBM_OfficeVision, simple databases like dBase [^15] https://www.dbase.com, and early file-sharing networks emerged. Bugs were often reported by users via email or phone calls to customer support. The software itself was usually shipped on floppy disks, and users would sometimes mail back faulty disks along with registration forms. When a customer support representative received a bug report, they would write the reproduction steps into a text file. This file would then be passed to a developer who would manually try to reproduce the bug. Fixing a bug involved manual code changes, which were then shared with the customer via a new floppy disk. Verification was done through manual testing with a documented checklist. The process was very slow and inefficient. While remote communication was a new development, collaboration was still low, often leading to a high TTR.
@@ 2.1.3. Internet Era (1980s-1990s)
This era saw the first dedicated bug tracking systems. Tools like GNATS (GNU Bug Tracking System) [^16] https://www.gnu.org/software/gnats/, Debbugs [^17] https://en.wikipedia.org/wiki/Debbugs, and CMVC (Yu, 1994) were introduced. GNATS, for instance, was an early, open source bug tracking system that used text-based files and email for communication, providing a structured interface for logging and tracking bugs. Debbugs was the software powering the Debian project’s issue tracking system. It did not have any form of web interface to edit bug reports, all modification was done through email. These systems allowed for detailed reproduction steps to be logged and were accessible to the whole team. When a bug was fixed, the code was committed to a version control system (VCS), and the bug report was updated with details about the fix.
@@
To manage the bug backlog, Netscape [^18] https://isp.netscape.com/ team leaders held weekly bug meetings, or ”bug councels” (Cusumano and Yoffie, 1999). As the project neared its release, these meetings increased in frequency to once or twice a day to prioritize and make decisions on fixes. These meetings were typically led by a project or release manager and included representatives from development, QA, and marketing teams. Netscape used a five-level classification system: critical, major, normal, minor, and trivial. Bugs were classified and entered into the database by the engineers who found them, typically from the QA team. The marketing team also maintained a ”top 10 bugs” list based on customer feedback, which was passed to QA.
@@ 2.1.4. Web-Based Era (2000s)
This decade saw a major shift to web-based platforms. Tools like Bugzilla, MantisBT [^19] https://mantisbt.org/, Trac [^20] https://trac.edgewall.org/, and early versions of Jira emerged. Bugzilla became a landmark open-source system. Bugzilla introduced foundational abstractions like structured fields (e.g., severity, priority, component), status transitions, and user roles. It provided a web-based interface and notification mechanisms, enabling better coordination among developers and testers (Mockus et al., 2002). While Bugzilla established the foundational blueprint for web-based tracking, Jira eventually surpassed it by offering comprehensive project management features and a flexible architecture that became the dominant choice for enterprise and agile teams. These systems treated bug reports as work items in a sprint, integrating with the emerging agile methodologies. The web UI made it much easier for everyone to access and track bugs. The expansion of community-contributed plug-ins and themes also reflected the diverse needs of development teams, from small startups to large open source foundations (Bertram et al., 2010). Collaboration improved with shared dashboards and email notifications, though bottlenecks were still common due to a lack of deep integration with development tools (Eren et al., 2023).
2.1.5. DevOps and Automation Era (2010s-2022)
In this era, platforms like GitHub Issues, GitLab Issues [^21] https://docs.gitlab.com/user/project/issues/, YouTrack [^22] https://www.jetbrains.com/youtrack/, and Azure DevOps [^23] https://azure.microsoft.com/en-us/products/devops have emerged as popular bug tracking systems. Bug tracking became fully integrated into the development lifecycle. GitHub Issues, for example, is part of a larger ecosystem that includes version control and continuous integration. Fixes were no longer just committed; they became a part of a Continuous Integration/Continuous Deployment (CI/CD) pipeline, where code changes automatically trigger tests and deployments.
The growing popularity of issue-driven development, where code commits are directly tied to issue identifiers, further solidified the role of bug tracking systems as central elements of the software delivery process. Developers began to view bug tracking systems not merely as databases or interfaces for bug database operations, but as vital components of planning, communication, and accountability (Gousios et al., 2014; Bird et al., 2009).